# Quantum geometry from the Moyal product: quantum kinetic equation and non-linear response

### v1

@@ 3.3 Fermi liquids with nontrivial band geometry
×$\quad\quad\times\left(\partial_{k_{l}}{\sf g}_{ij}(k)q^{\prime\prime}_{i}q^{\prime\prime}_{j}q^{\prime}_{l}+i{\sf\Omega}_{ij}(k)q^{\prime\prime}_{i}q^{\prime}_{j}\right)+\left(\frac{(q^{\prime}_{i}+q^{\prime\prime}_{i})q^{\prime}_{j}q^{\prime\prime}_{l}}{\omega^{\prime\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime\prime}}-\frac{q^{\prime}_{i}q^{\prime\prime}_{j}q^{\prime}_{l}}{\omega^{\prime}+\omega^{\prime\prime}-\vec{v}\cdot(\vec{q}^{\prime}+\vec{q}^{\prime\prime})}\right)2\hbar v_{l}{\sf t}_{ij}(k)\Bigg{]},$
@@
where $k_{\mathrm{F}}$ is the Fermi wavevector and ${\sf g}_{ij,\mathrm{F}}(\theta)={\sf g}_{ij}(k=k_{\mathrm{F}},\theta)$. Notice that, as we assumed $\omega\ll\omega_{\mathrm{gap}}$, (128) only involves intraband contributions. In addition to the leading term, $\propto q$ ($\frac{\hbar^{2}k_{\rm F}}{2\pi^{2}}|\vec{q}|$), which is the single-band contribution to the Fermi gas, the quantum metric contributes at cubic order, $q^{3}$, to $S_{q}^{\mathrm{intra}}$ provided the $\theta$-integral in (128) is non-vanishing. These results can be compared to a Green’s function calculation of non-interacting multi-band fermions, which we carry out in appendix L.
@@ 4.1 Framework and comparison to other approaches
In this paper, we have begun with the full matrix quantum kinetic equation and systematically derived a band-diagonal one to second order in the semi-classical expansion, which can be regarded as a formal expansion in $\hbar$, but is practically controlled as an expansion in spatial gradients. This is one order beyond the standard approach—first order leads to Berry curvature effects but not those of the quantum metric. [^4] One may wonder why they appear at different orders even if they can be united in a QGT. This is because at the leading order, the Berry curvature is linear in the Berry connection but the quantum metric is not, which itself comes from cancellations due to the fact that the Moyal product is antisymmetric but the quantum metric is symmetric. The diagonal dynamics enjoys a reduction of degrees of freedom: for an $N$-band system, it is expressed in terms of $N$ phase space densities rather than $N^{2}$ matrix elements. While this offers some computational gain, the main advantage is conceptual: one sees directly how all physics can be systematically expressed in terms of band-diagonal quantities, with quantum geometry emerging naturally through the diagonalization.
A recent preprint by Mitscherling et al. takes a similar approach in which they define a matrix-valued Wigner distribution function to study transport [34]. While their work is similar in spirit to ours, their formalism is developed for lattices and they do not include the effects of band geometry.
@@ 4.3 Ramifications of the gradient expansion
The second ambiguity is more treacherous. The (transport) current itself can be shifted by any curl, $J_{i}\rightarrow J_{i}+\partial_{x_{j}}M_{ij}$, where $M_{ij}$ is an anti-symmetric “magnetization density”, even when the definition of the density $\rho$ is unchanged. Due to the spatial derivative, this is not expected to change the zero momentum current in a translationally invariant system (for which Fourier decomposition is sensible), but it does affect non-zero wavevector currents, and it does affect the local currents when the system is not translationally invariant. The latter situation is important, for example, in the presence of even a constant temperature gradient, which leads to subtleties in calculations of the thermal Hall effect. However, by construction, the magnetization currents do not contribute to net transport through any closed surface or a surface that terminates outside the sample. This means that while they can, and do. affect calculations of the conductivity, the magnetization currents do not affect the conductance. In the formalism of this paper, it is possible to calculate a true conductance by treating a finite sample and obtaining the integrated current across a cross-section. We emphasize that the “Biot-Savart current” density, which is defined as the one that appears in Maxwell’s equations, is unambiguous. This is the one we derived in section 2.8, and has been used in this work.
@@ 4.5 Extensions
There are a number of interesting directions to extend this work. One could apply the formalism developed here to study thermal and thermo-electric currents and responses. It is particularly well-suited for this because a spatially inhomogeneous temperature can be explicitly treated (as can a non-thermal non-equilibrium state) without recourse to the “trick” of inserting an artificial gravitational field à la Luttinger, whose applicability beyond linear response is questionable. It would be natural to extend the current formalism to superonducting states, i.e. Bogoliubov-de Gennes Hamiltonians. An important direction is to include the effects of electron-electron or electron-phonon interactions. This could be accomplished via the Keldysh method, in which the semi-classical expansion is well-established for single band systems and can be best carried out by extending the Wigner transformation from just the space/momentum domain to include as well the time/frequency variables. The latter is essential to take into account frequency-dependent self-energy effects. We believe the techniques developed here should be a foundation for such a study in multi-band Fermi liquids. Another interesting direction is to extend our formalism to consider finite frequency responses beyond the low-frequency limit we considered in this work. This would involve solving for the off-diagonal components of the distribution function using eq. 32. This would allow us to study photovoltaic effects, shift and injection currents, and sum rules of optical conductivity which have been shown to be related to the integrated quantum metric [35, 37, 38]. Lastly, our formalism was developed assuming non-degenerate energy states but it would be of interest to extend our formalism to be fully compatible with degeneracies which generically occur in crystalline systems.
@@ Appendix C Defining $f,h$
$f_{n}\equiv\mathrm{tr}[U\star p_{n}\tilde{f}_{n}\star U^{\dagger}].$ (134)
@@ Appendix D Diagonalization by flow equations
$\begin{split}\hbar\partial_{\hbar}\tilde{h}=&\tilde{h}^{(1)}+2\tilde{h}^{(2)}+\cdots=[\tilde{h}\,\overset{\star}{,}\,Y]+\frac{\hbar}{2}\omega^{\alpha\beta}\{\Lambda_{\alpha}\,\overset{\star}{,}\,\partial_{\beta}\tilde{h}\}+\frac{i\hbar}{4}\omega^{\alpha\beta}\{\Lambda_{\alpha}\,\overset{\star}{,}\,[\Lambda_{\beta}\,\overset{\star}{,}\,\tilde{h}]\}.\end{split}$ (143)
@@ Appendix E Derivation of the equilibrium distribution
where $n_{F}^{(k-1)}$ is the $k-1$-th derivative of $n_{F}(x)=(e^{\beta(x-\mu)}+1)^{-1}$ which is the Fermi-Dirac distribution function. We then find
@@
$\tilde{f}_{\mathrm{eq}}=n_{\mathrm{F}}(-\mu)+n_{\mathrm{F}}^{\prime}(-\mu)\tilde{h}+\frac{1}{2}n_{\mathrm{F}}^{\prime\prime}(-\mu)\tilde{h}\star\tilde{h}+\frac{1}{3!}n_{\mathrm{F}}^{\prime\prime\prime}(-\mu)\tilde{h}\star\tilde{h}\star\tilde{h}+\ldots,$ (170)
which reduces to $n_{\mathrm{F}}(h-\mu)$ if all the star products are the normal products. By properly arranging the infinite series, we find it is precisely (169) to the second order in $\hbar$.
Appendix F Derivation of the electric current eq. 46
@@
We first prove eq. 173 by considering a separable Hamiltonian meaning $\hat{H}=H_{0}(\hat{p})+H_{1}(\hat{x})$. Then, we consider the more general case.
We assume spinless fermions and introduce the external field via minimal coupling. Then, $\hat{H}_{\mathcal{A}}=H_{0}(\hat{p}+e\mathcal{A}(\hat{x}))+H_{1}(\hat{x})$. The second term $H_{1}(\hat{x})$ does not contribute to the current, so we ignore it. In general $H_{0}(p)$ can be expressed as
@@
where $G_{\alpha_{1}\cdots\alpha_{n}}$ is a matrix-valued symmetric tensor. The external gauge field can be introduced by replacing each $\hat{p}_{\alpha}$ with $\hat{\pi}_{\alpha}=\hat{p}_{\alpha}+e\mathcal{A}(\hat{x})$. Applying the Wigner transformation we get
@@
$H_{0}(\hat{p})\rightarrow$ $\,{\sf H}_{0}(p)=\sum_{n=1}^{\infty}\frac{1}{n!}G_{\alpha_{1}\cdots\alpha_{n}}\,p_{\alpha_{1}}\cdots p_{\alpha_{n}}$ (177)
$H_{0}(\hat{\pi})\rightarrow$ $\,{\sf H}_{0\mathcal{A}}(x,p)=\sum_{n=1}^{\infty}\frac{1}{n!}G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}$ (178)
@@
{mdframed}
Lemma
$G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $G_{\alpha_{1}\cdots\alpha_{n}}\left(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}-\frac{e\hbar^{2}}{24}n(n-1)(n-2)\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}}\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}(x)}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$
$=$ $\left(1-\frac{e\hbar^{2}}{24}\frac{\partial^{2}\mathcal{A}_{\mu}(x)}{\partial x_{\nu}\partial x_{\gamma}}\frac{\partial^{3}}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}\right)G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}$ (179)
Proof: First, note $\left\{A\,\overset{\star}{,}\,B\right\}/2=AB-(\hbar^{2}/8)\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}A\partial^{2}_{\sigma\lambda}B+\mathcal{O}(\hbar^{3})$. Then, using the fact that $G_{\alpha_{1}\cdots\alpha_{n}}$ is symmetric under exchange of its indices we can get
$G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $\frac{1}{2}G_{\alpha_{1}\cdots\alpha_{n}}\,\left\{\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}}\,\overset{\star}{,}\,\pi_{\alpha_{1}}\right\}$
$=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left((\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}})\pi_{\alpha_{1}}-\frac{\hbar^{2}}{8}\frac{\partial^{2}}{\partial p_{\nu}\partial p_{\gamma}}(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n-1}})\frac{\partial^{2}\pi_{\alpha_{n}}}{\partial x_{\nu}\partial x_{\gamma}}\right)+\mathcal{O}(\hbar^{3})$
$=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left((\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}})\pi_{\alpha_{1}}-(n-1)(n-2)\frac{e\hbar^{2}}{8}(\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}})\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$
Notice this sets up a recursive relation. The first term in the last line with the $\star$ products is the left-hand side up to differences in the symmetric tensor. If we continue this recursive relation we find
$G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}-\sum_{m=1}^{n-2}\frac{(n-m)!}{(n-m-2)!}\frac{e\hbar^{2}}{8}(\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}})\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$
Finally, we use the fact that $\sum_{m=1}^{n-2}\frac{(n-m)!}{(n-m-2)!}=\frac{1}{3}n(n-1)(n-2)$ and we get (179). Applying the lemma (179) we find
@@
where $\{v_{n}\},\{k_{n}\}$ are a set of polynomial functions and $\Gamma_{n}$ is a constant matrix. We now once again couple the system to an external gauge field by taking $\hat{p}\rightarrow\hat{\pi}$, and calculate the Wigner transformation. We know what the Wigner transformation of $k_{n}(\hat{\pi})$ must be from (180). Therefore,
@@
Appendix G Projection operator identities

### v2

@@ 3.3 Fermi liquids with nontrivial band geometry
$\quad\quad\times\left(\partial_{k_{l}}{\sf g}_{ij}(k)q^{\prime\prime}_{i}q^{\prime\prime}_{j}q^{\prime}_{l}+i{\sf\Omega}_{ij}(k)q^{\prime\prime}_{i}q^{\prime}_{j}\right)+\left(\frac{(q^{\prime}_{i}+q^{\prime\prime}_{i})q^{\prime}_{j}q^{\prime\prime}_{l}}{\omega^{\prime\prime}-\vec{v}\cdot{\vec{q}\,}^{\prime\prime}}-\frac{q^{\prime}_{i}q^{\prime\prime}_{j}q^{\prime}_{l}}{\omega^{\prime}+\omega^{\prime\prime}-\vec{v}\cdot(\vec{q}^{\prime}+\vec{q}^{\prime\prime})}\right)2\hbar v_{l}{\sf t}_{ij}(k)\Bigg],$
@@
where $k_{\mathrm{F}}$ is the Fermi wavevector and ${\sf g}_{ij,\mathrm{F}}(\theta)={\sf g}_{ij}(k=k_{\mathrm{F}},\theta)$. Notice that, as we assumed $\omega\ll\omega_{\mathrm{gap}}$, (128) only involves intraband contributions. In addition to the leading term, $\propto q$ ($\frac{\hbar^{2}k_{\rm F}}{2\pi^{2}}|\vec{q}|$), which is the single-band contribution to the Fermi gas, the quantum metric contributes at cubic order, $q^{3}$, to $S_{q}^{\mathrm{intra}}$ provided the $\theta$-integral in (128) is non-vanishing. These results can be compared to a Green’s function calculation of non-interacting multi-band fermions, which we carry out in appendix M.
@@ 4.1 Framework and comparison to other approaches
In this paper, we have begun with the full matrix quantum kinetic equation and systematically derived a band-diagonal one to second order in the semi-classical expansion, which can be regarded as a formal expansion in $\hbar$, but is practically controlled as an expansion in spatial gradients. This is one order beyond the standard approach—first order leads to Berry curvature effects but not those of the quantum metric. [^5] One may wonder why they appear at different orders even if they can be united in a QGT. This is because at the leading order, the Berry curvature is linear in the Berry connection but the quantum metric is not, which itself comes from cancellations due to the fact that the Moyal product is antisymmetric but the quantum metric is symmetric. The diagonal dynamics enjoys a reduction of degrees of freedom: for an $N$-band system, it is expressed in terms of $N$ phase space densities rather than $N^{2}$ matrix elements. While this offers some computational gain, the main advantage is conceptual: one sees directly how all physics can be systematically expressed in terms of band-diagonal quantities, with quantum geometry emerging naturally through the diagonalization. Our results can be seen as an extension of the pioneering approach of Blount[2, 3, 4], which also considered Moyal diagonalization, but to lower order and with less generality (and our presentation is significantly more self-contained than Blount’s). A more detailed comparison to some of Blount’s work in found in Appendix N.
A recent preprint by Mitscherling et al. takes an approach in which they define a matrix-valued Wigner distribution function to study transport [19]. Their formalism which is developed for lattice systems, involves an expansion in spatial gradients which corresponds to first-order in $\hbar$ of our results.
@@ 4.3 Ramifications of the gradient expansion
The second ambiguity is more treacherous. The (transport) current itself can be shifted by any curl, $J_{i}\rightarrow J_{i}+\partial_{x_{j}}M_{ij}$, where $M_{ij}$ is an anti-symmetric “magnetization density”, even when the definition of the density $\rho$ is unchanged. Due to the spatial derivative, this is not expected to change the zero momentum current in a translationally invariant system (for which Fourier decomposition is sensible), but it does affect non-zero wavevector currents, and it does affect the local currents when the system is not translationally invariant. The latter situation is important, for example, in the presence of even a constant temperature gradient, which leads to subtleties in calculations of the thermal Hall effect. However, by construction, the magnetization currents do not contribute to net transport through any closed surface or a surface that terminates outside the sample. This means that while they can and do affect calculations of the conductivity, the magnetization currents do not affect the conductance. In the formalism of this paper, it is possible to calculate a true conductance by treating a finite sample and obtaining the integrated current across a cross-section. We emphasize that the “Biot-Savart current” density, which is defined as the one that appears in Maxwell’s equations, is unambiguous. This is the one we derived in section 2.8, and has been used in this work  [^7] Although we do not attempt to distinguish the transport and magnetization current in this work, they may be separated by considering the Joule heating, c.f. Ref. [34]. .
@@ 4.5 Extensions
There are a number of interesting directions to extend this work. One could apply the formalism developed here to study thermal and thermo-electric currents and responses. It is particularly well-suited for this because a spatially inhomogeneous temperature can be explicitly treated (as can a non-thermal non-equilibrium state) without recourse to the “trick” of inserting an artificial gravitational field à la Luttinger, whose applicability beyond linear response is questionable. It would be natural to extend the current formalism to superonducting states, i.e. Bogoliubov-de Gennes Hamiltonians. An important direction is to include the effects of electron-electron or electron-phonon interactions. This could be accomplished via the Keldysh method, in which the semi-classical expansion is well-established for single band systems and can be best carried out by extending the Wigner transformation from just the space/momentum domain to include as well the time/frequency variables. The latter is essential to take into account frequency-dependent self-energy effects. We believe the techniques developed here should be a foundation for such a study in multi-band Fermi liquids. Another interesting direction is to extend our formalism to consider finite frequency responses beyond the low-frequency limit we considered in this work. This would involve solving for the off-diagonal components of the distribution function using eq. 32. This would allow us to study photovoltaic effects, shift and injection currents, and sum rules of optical conductivity which have been shown to be related to the integrated quantum metric [35, 13, 37, 38]. Lastly, our formalism was developed assuming non-degenerate energy states but it would be of interest to extend our formalism to be fully compatible with degeneracies which generically occur in crystalline systems.
@@ Appendix C Defining $f,h$
$f_{n}\equiv\mathrm{tr}[U\star p_{n}\tilde{f}_{n}\star U^{\dagger}]=\mathrm{tr}[U\star p_{n}\tilde{f}\star U^{\dagger}].$ (134)
Note that as a consequence of this definition, when $\tilde{\mathcal{F}}=0$, one has
$\mathrm{tr}[\mathsf{F}]=\mathrm{tr}[f],\qquad\tilde{\mathcal{F}}=0,$ (135)
so that the phase space density is given by $\mathrm{tr}[f]$ in this situation (note that here the identity holds without a phase space integral).
@@ Appendix D Diagonalization by flow equations
[…]
@@ Appendix E Derivation of the equilibrium distribution
where $n_{F}^{(k-1)}$ is the k1k-1-th derivative of $n_{F}(x)=(e^{\beta(x-\mu)}+1)^{-1}$ which is the Fermi-Dirac distribution function. We then find
@@
$\tilde{f}_{\mathrm{eq}}=n_{\mathrm{F}}(-\mu)+n_{\mathrm{F}}^{\prime}(-\mu)\tilde{h}+\frac{1}{2}n_{\mathrm{F}}^{\prime\prime}(-\mu)\tilde{h}\star\tilde{h}+\frac{1}{3!}n_{\mathrm{F}}^{\prime\prime\prime}(-\mu)\tilde{h}\star\tilde{h}\star\tilde{h}$$\ldots$, (171)
which reduces to $n_{\mathrm{F}}(h-\mu)$ if all the star products are the normal products. By properly arranging the infinite series, we find it is precisely (170) to the second order in $\hbar$; this agrees with a result of Blount[2].
Appendix F Polarization in an insulator
Let us assume an insulator so that $n_{\mathrm{F}}=1$ throughout the Brillouin zone. From (38), we find the bound charge density is given by
$\rho_{b}\equiv-e\int_{p}(f_{\mathrm{eq}}-n_{\mathrm{F}})=-e\hbar\int_{p}\left(\Omega_{x^{i}p^{i}}+\frac{\hbar}{2}\left(\Omega_{x^{i}x^{j}}\Omega_{p^{i}p^{j}}-\Omega_{x^{i}p^{j}}\Omega_{p^{i}x^{j}}-\Omega_{x^{i}p^{i}}\Omega_{x^{j}p^{j}}+\partial_{x^{i}x^{j}}^{2}g_{p^{i}p^{j}}\right)\right)+\mathcal{O}(\hbar^{3}).$ (172)
The polarization vector $\mathcal{P}_{i}$ is defined through $\rho_{b}=-\partial_{i}\mathcal{P}_{i}$. First, the single Berry curvature term in (172) gives
$\mathcal{P}_{i}^{(1)}=e\hbar\int_{p}\left(A_{p^{i}}+\hbar\left(A_{p^{i}}\partial_{x^{j}}A_{p^{j}}+A_{x^{j}}\partial_{p^{j}}A_{p^{i}}+A_{p^{j}}\partial_{p^{i}}A_{x^{j}}\right)\right)+\mathcal{O}(\hbar^{3}).$ (173)
Second, the double Berry curvature terms give
$\mathcal{P}_{i}^{(2)}=\frac{e\hbar^{2}}{2}\int_{p}\left(A_{p^{j}}\partial_{x^{j}}A_{p^{i}}-A_{p^{i}}\partial_{x^{j}}A_{p^{j}}-2A_{p^{j}}\partial_{p^{i}}A_{x^{j}}-2A_{x^{j}}\partial_{p^{j}}A_{p^{i}}\right)+\mathcal{O}(\hbar^{3}).$ (174)
Then, the quantum metric gives
$\mathcal{P}_{i}^{(3)}=\frac{e\hbar^{2}}{2}\int_{p}\partial_{x^{j}}g_{p^{i}p^{j}}+\mathcal{O}(\hbar^{3}).$ (175)
Gathering above, we find the total polarization vector as
$\mathcal{P}_{i}=e\hbar\int_{p}\left(A_{p^{i}}+\frac{\hbar}{2}\partial_{x^{j}}\left(A_{p^{i}}A_{p^{j}}\right)+\frac{\hbar}{2}\partial_{x^{j}}g_{p^{i}p^{j}}\right)+\mathcal{O}(\hbar^{3}).$ (176)
Comparing with the semiclassical result in [42], we find that only the contribution in (173) can be identified the latter, in particular with Eq.(1), Eq.(3) and Eq.(11) in [42]. The remaining terms from $\mathcal{P}_{i}^{(2)}$ and $\mathcal{P}_{i}^{(3)}$ are distinct, but are not present in [42]. Nevertheless, our final result passes all physical checks and the discrepancy is an interesting subject for further study. In our opinion our calculation is more rigorous.
For a separable Hamiltonian (c.f. (89)), the gauge field $A_{p^{i}}$ and the quantum metric $g_{p^{i}p^{j}}$ are functions of only $p$ at the leading order, so the spatial derivatives vanish, and the resulting polarization $\mathcal{P}_{i}=e\hbar\int_{p}A_{p^{i}}+\mathcal{O}(\hbar^{3})$ involves only the second order corrections to the gauge field $A_{p^{i}}$, recovering the semiclassical results [9].
Appendix G Derivation of the electric current eq. 46
@@ Appendix F Polarization in an insulator
[…]
@@
[…]
@@
[…]
@@
Lemma $G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $G_{\alpha_{1}\cdots\alpha_{n}}\left(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}-\frac{e\hbar^{2}}{24}n(n-1)(n-2)\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}}\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}(x)}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$ $=$ $\left(1-\frac{e\hbar^{2}}{24}\frac{\partial^{2}\mathcal{A}_{\mu}(x)}{\partial x_{\nu}\partial x_{\gamma}}\frac{\partial^{3}}{\partial\pi_{\mu}\partial\pi_{\nu}\partial\pi_{\gamma}}\right)G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}$ (185) Proof: First, note $\left\{A\,\overset{\star}{,}\,B\right\}/2=AB-(\hbar^{2}/8)\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}A\partial^{2}_{\sigma\lambda}B+\mathcal{O}(\hbar^{3})$. Then, using the fact that $G_{\alpha_{1}\cdots\alpha_{n}}$ is symmetric under exchange of its indices we can get $G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $\frac{1}{2}G_{\alpha_{1}\cdots\alpha_{n}}\,\left\{\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}}\,\overset{\star}{,}\,\pi_{\alpha_{1}}\right\}$ $=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left((\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}})\pi_{\alpha_{1}}-\frac{\hbar^{2}}{8}\frac{\partial^{2}}{\partial p_{\nu}\partial p_{\gamma}}(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n-1}})\frac{\partial^{2}\pi_{\alpha_{n}}}{\partial x_{\nu}\partial x_{\gamma}}\right)+\mathcal{O}(\hbar^{3})$ $=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left((\pi_{\alpha_{2}}\star\cdots\star\pi_{\alpha_{n}})\pi_{\alpha_{1}}-(n-1)(n-2)\frac{e\hbar^{2}}{8}(\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}})\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$ Notice this sets up a recursive relation. The first term in the last line with the $\star$ products is the left-hand side up to differences in the symmetric tensor. If we continue this recursive relation we find $G_{\alpha_{1}\cdots\alpha_{n}}\,\pi_{\alpha_{1}}\star\cdots\star\pi_{\alpha_{n}}=$ $G_{\alpha_{1}\cdots\alpha_{n}}\,\left(\pi_{\alpha_{1}}\cdots\pi_{\alpha_{n}}-\sum_{m=1}^{n-2}\frac{(n-m)!}{(n-m-2)!}\frac{e\hbar^{2}}{8}(\pi_{\alpha_{4}}\cdots\pi_{\alpha_{n}})\frac{\partial^{2}\mathcal{A}_{\alpha_{1}}}{\partial x_{\alpha_{2}}\partial x_{\alpha_{3}}}\right)+\mathcal{O}(\hbar^{3})$ Finally, we use the fact that $\sum_{m=1}^{n-2}\frac{(n-m)!}{(n-m-2)!}=\frac{1}{3}n(n-1)(n-2)$ and we get (185).
Applying the lemma (185) we find
@@
[…]
@@
Appendix H Projection operator identities